# EvoChart: A Benchmark and a Self-Training Approach Towards Real-World Chart Understanding

### v1

@@ 1 Introduction
Chart Question Answering (CQA) aims to answer specific questions based on the context provided by chart images, enabling automated data analysis, such as the business data reports. This process requires complex chart understanding and visual reasoning skills to interpret various elements, including visual components, text and values. Consequently, CQA tasks have attracted the interest of researchers [10, 24, 20].
Recently, VLMs [18, 7, 15] have shown significant advancements in general visual capabilities, especially in chart understanding, achieving high scores on the ChartQA dataset [35, 4, 23]. However, their real-world performance is notably weaker than their ChartQA test set performance. We conducted a test to illustrate this, as shown in Figure 1, we discarded complex reasoning problems in the ChartQA [20] training set and posed 103 basic understanding questions. We then evaluated various VLMs on these questions. The results, presented in the Appendix, show that performance dropped by over 40% compared to ChartQA scores, even for questions on training set charts. This highlights two points: first, current VLMs are capable of answering some chart-reasoning questions, but they lack a comprehensive understanding of charts. Second, the ChartQA dataset allows models to correctly answer questions without a comprehensive understanding of the charts, leading to an overestimation of the capabilities of current models [30].
@@
Firstly, the lack of high-quality chart training data is a major reason why current models lack robust chart understanding capabilities. Existing methods [20, 19, 23] for collecting chart training data fall into two categories: manual annotation and automatic synthesis. Manually annotated data have real-world chart appearances but suffer from coarse granularity and high human costs. Automatically synthesized data offer fine-grained annotations but lack real-world diversity, leading to poor model robustness. Thus, constructing chart datasets is a challenging balance between cost and quality, resulting in a scarcity of high-quality chart training data.
@@
  • We conducted extensive experiments on the EvoChart method and EvoChart-QA. Results demonstrate that the EvoChart method significantly outperforms other data synthesis methods, and we also deeply analyze the performance of various VLMs on EvoChart-QA.We have made EvoChart and EvoChart-QA publicly available at https://github.com/MuyeHuang/EvoChart
@@ 2.1 Chart Question Answering Datasets
Since FigureQA [12] pioneered the CQA task, numerous datasets for chart question answering have emerged. Synthetic datasets, such as DVQA [10], PlotQA [24], RealCQA [2], ChartX [31] and UniChart [19]. This datasets utilize synthetically generated charts or templat-base questions. Generate datasets such as ChartSFT [23], utilize a mixture of GPT-4 [25]-generated charts and questions. Mixed datasets as ChartQA [20] and Charxiv [30], the former is a dataset compiled semi-manually with the assistance of templates, while the latter is template-based and requires evaluation by GPT-4o. In contrast, EvoChart-QA focus on real-world scenarios and employ an automated evaluation method that does not necessitate the utilization of GPT-4.
@@ 2.2 Visual Language Models on CQA
VLMs are language models with visual understanding capabilities, and they have numerous applications in CQA tasks. Small VLMs like ChartReader [6], MatCha [16] and UniChart [19] have shown superior performance on tasks like PlotQA and DVQA, highlighting the potential of VLMs in CQA. ChartLlama [9] was a milestone, being the first to apply LLaVa1.5 [17] to CQA tasks and achieving impressive performance. Subsequently, works like ChartInstruct [21], ChartAst-D [23], and TinyChart [35] delved into the multimodal alignment and CQA reasoning aspects of VLMs in CQA, achieving remarkable performance. Recently, open-source general VLMs such as Phi3-Vision [1] and Intern-VL2.0 [5], through large-scale training, have achieved state-of-the-art performance on the ChartQA dataset.
@@ 2.3 Self-Training Approach
With the increasing capabilities of language models, numerous researchers have begun to explore the potential of leveraging language models for self-training. GPT3Mix [34] proved that large language model augmentation of textual corpora is very effective. Further, ReST [8] achieves cost-effective and efficient human preference alignment through a dual-loop self-training approach. Dennis et al. [28] obtained new data from the self-talk of multi-role-playing LLM Agents by adding a filtering check mechanism, realizing efficient self-training. Recently, Xu et al. [32] proposed ENVISIONS, which uses a neural-symbolic self-training approach to significantly improve mathematical and logical reasoning abilities without relying on external stronger models or evaluation tools. Inspired by their work, our proposed EvoChart focuses on a scalable self-training process.
@@ 3.1 Compositional Chart Generation
Compositional chart generation aims to produce high-quality and diverse charts at minimal cost, serving as the core component of EvoChart’s construction. Previous approaches [9, 23, 35] either rely solely on GPT-4 for chart generation result in limited diversity and high costs, or using plotting libraries for random generation often leads to unrealistic themes and styles. To achieve continuous, low-cost, and diverse chart creation, we propose a two-step generation strategy within the compositional chart generation process:
1) Chart Code Seed Generation: This step serves as the initial phase of chart generation, aiming to produce fundamental chart code containing elements that are difficult to achieve via random processes. These elements include chart themes and appropriate units for the x- and y-axes. Notably, this step is executed only once during the entire chart dataset construction process. As [33] demonstrated the feasibility of large language models for code generation, we employ sophisticated prompt engineering techniques to guide GPT-4 in generating a large number of real-world chart code seeds. Our prompts for GPT-4 are based on the following key aspects:
@@
Through these cost-effective, one-time steps, we have generated 25,000 chart code seeds. These seeds exhibit diverse themes and color palettes, and retain hundreds of rich configuration options, forming the foundation for subsequent low-cost data generation.
@@
Numeric Labels: The numeric label within charts is a crucial configurable option. For each legend in chart, we randomly decide whether to display numeric labels. If labels are displayed, their positions relative to the data points (above, below, etc.) are also randomly assigned. For label font color, we randomly choose whether it should match the legend color or be assigned a different color independently.
@@
Other Configurations: A multitude of detailed configurations influence chart appearance, including line width, line style (solid or dashed), bar stacking, axis visibility, font size, font type, and more. We introduce randomness into these configurations through a range of selectable options, and we employ ECharts [13] for rendering charts.
@@ 3.2 Chart Evaluation and Refinement
Stage Type Line Bar Pie Scatter Total
Stage1 Chart 9,074 9,772 4,594 3,878 27,318
Query 315,902 203,898 31,533 90,851 642,184
Stage2 Chart 17,280 17,293 7,720 6,259 48,462
Query 599,078 360,827 52,866 146,631 1,159,402
Stage3 Chart 24,279 22,562 11,004 9,316 67,161
Query 841,745 470,771 75,356 218,227 1,606,099
Table 1: EvoChart Corpus statistics
3.3 QA-pair Generation and Training
QA-pair Generation and Training aims to generate question-answer pairs based on charts, incorporating these data into the EvoChart Corpus and training the EvoChart model for the next stage. We generate question-answer pairs using various question templates. Notably, we focus on basic chart understanding, the templates specifically focus on the alignment of visual-text-value information in charts (e.g., extracting values through visual information, extracting visual information through text). Additionally, we generate rich question-CoT pairs using composable vCoTs [26] and distinguish Direct from vCoT using Instruct. Since vCoTs solely serve to enhance model comprehension, we only mix in 20% of vCoT data in the training data. We have established 198 question templates with corresponding answers including Direct and over 500 vCoT templates. The statistical data for the EvoChart Corpus is presented in Table 1. A detailed template can be found in the Appendix.
@@ 4 EvoChart-QA Benchmark
4.1 Chart Selection
To enable EvoChart-QA to emulate real-world chart understanding scenarios, all charts in our dataset are carefully selected by human experts. Specifically, we crawled 1,000 charts from 140 different websites. Human experts then filtered out images with ambiguous meanings or damages, resulting in a final dataset of 625 valid images. These images include line charts, bar charts, pie charts, and scatter plots. The distribution of these images is shown in Table 2. Examples and sources of all images are detailed in the Appendix.
Dataset Type Line Bar Pie Scatter Total
ChartQA-test Chart 134 405 86 - 625
Query 268 810 172 - 1,250
EvoChart-QA Chart 200 200 125 125 650
Query 400 400 225 225 1,250
Table 2: Comparison of EvoChart-QA and ChartQA Benchmark
4.2 Question Construction
We focus on chart basic understanding questions. Following the definitions of prior researchers [11, 24], we concentrate on data and structural retrieval questions. Specifically, we categorize the problems into two types during manual construction: Direct Retrieval and Complex Retrieval. Direct Retrieval questions focus on understanding the image and directly extracting its content, while Complex Retrieval questions emphasize performing multiple visual reasoning steps on the chart.
1) Direct Retrieval. Direct Retrieval aims to directly extract elements from a chart based on the question. To comprehensively assess the model’s ability to extract various elements from charts, we categorize chart elements into three types: label, value, and visual. Label elements refer to textual content in the chart, such as chart title, axis labels, etc. Value elements refer to data values conveyed by the chart, which are displayed or implicitly provided based on the chart author’s intention. Visual elements refer to all visual descriptions in the chart, such as line color, largest segment, etc. For example: “What is the value of the green dashed line in 2015?” Although Direct Retrieval only involves extracting elements from charts, it remains a challenging task. On the one hand, real-world charts often exhibit non-standard variations. For example, they may include rich text such as numerous logos inserted in the image, or they may combine multiple chart forms within a single chart to facilitate expressions. On the other hand, questions posed by real-world users may contain ambiguous expressions. For example, “Which country’s total GDP is represented by the bar in the lightest shade of blue?” This visual description is ambiguous, but it is a clearly question for human observers.
@@ 5.1 Setup
Datasets. To comprehensively evaluate the effectiveness of EvoChart, we chose to test it on both ChartQA and EvoChart-QA. ChartQA [20] is a dataset with two subsets: “Augment”, which is machine-generated, and “Human”, which is manually curated. “Augment” focuses on element extraction tasks within machine-synthesized images, while “Human” emphasizes complex numerical and logical reasoning tasks in real-world charts. EvoChart-QA is a novel real-world benchmark that we proposed.
Models. We conducted extensive evaluations on both open-source and proprietary models. For open-source models, we tested Phi3-Vision-4B [1], QwenVL-Chat-7B [3], LlaVa1.6-Vicuna-7B [17], Intern-VL-2.0-8B [5], Llama3-Llava-Next-8B [14], CogVLM2-19B [29], LlaVa1.6-YI-34B, Intern-VL-2.0-40B, ChartLlama-13B [9], ChartAst-S-13B [23], ChartIns-Llama2-7B [21], ChartIns-FlanT5-3B, ChartGemma-2B [22], and TinyChart-3B [35]. For proprietary models, we tested Gemini-1.5-Flash [27], Gemini-1.5-Pro, Qwen-VL-Plus, Qwen-VL-Max, GPT-4-turbo [25], and GPT-4o. For all models, we employed a zero-shot approach. The specific configurations of all models are provided in the Appendix.
Settings. In EvoChart method, we utilize Phi3-Vision [1] as the initialization model. We conducted a 3-Stage data synthesis and training process, with each Stage undergoing 1 Epoch of SFT with a learning rate of 2e-5 and using cosine learning rate scheduler. All experiments were completed on 4 NVIDIA A800 80G GPUs.
@@ 5.2 Experimental Results
Tables 4 and 5 present the performance of EvoChart and other open-source or proprietary VLMs on EvoChart-QA and ChartQA. We elaborate on the experimental results in two aspects: EvoChart-QA and EvoChart.
@@
2) All models demonstrate significantly weaker performance on Complex Retrieval compared to Direct Retrieval, indicating that reasoning over visual information poses a substantially greater challenge than direct extraction of information from charts. Furthermore, nearly all models exhibit lower accuracy on Pie and Scatter chart types compared to their average performance. This suggests that Pie and Scatter charts pose a greater challenge for these models to understand.
3) Open-source general-purpose models exhibit comparable chart comprehension abilities to proprietary models. This suggests that models pretrained on large-scale chart data possess strong generalization capabilities. However, while chart expert models fine-tuned on specific domains achieve impressive scores on the ChartQA dataset, their performance degrades significantly to below 30% when confronted with the entirely out-of-distribution (OOD) EvoChart-QA dataset.
@@
1) Although the EvoChart model is trained on synthetic data, it achieves state-of-the-art performance on the entirely OOD real-world benchmark EvoChart-QA and exhibits competitive performance on the chart reasoning task ChartQA. This validates the strong generalization ability of the EvoChart approach.
@@ 5.3 Analysis
EvoChart Ablation Study. We conducted comprehensive ablation studies on EvoChart, and the results are summarized in Table 6. We trained and generated EvoChart for 1 to 3 stages. Meanwhile, to verify the effectiveness of Chart Evaluation and Refinement, we set up EvoChart without refinement and trained it for 3 stages, denoted as “w/o refine stage-3.” Based on the experimental results presented in the Table6 and Tabel 1, we observed the following:
1) As the number of EvoChart method Stages increases, the scale of the EvoChart-Dataset expands, and the performance of the EvoChart steadily improves. This demonstrates the effectiveness of EvoChart, a multi-round self-training approach.

### v2

@@ 1 Introduction
Chart Question Answering (CQA) aims to answer specific questions based on the context provided by chart images, enabling automated data analysis, such as the business data reports. This process requires complex chart understanding and visual reasoning skills to interpret various elements, including visual components, text and values. Consequently, CQA tasks have attracted the interest of researchers (Kafle et al. 2018; Methani et al. 2020; Masry et al. 2022).
Recently, VLMs (Liu et al. 2023b; Dai et al. 2023; Lin et al. 2023; Zhu et al. 2024) have shown significant advancements in general visual capabilities, especially in chart understanding, achieving high scores on the ChartQA dataset (Zhang et al. 2024; Chen et al. 2023a; Meng et al. 2024). However, their real-world performance is notably weaker than their ChartQA test set performance. We conducted a test to illustrate this, as shown in Figure 1, we discarded complex reasoning problems in the ChartQA (Masry et al. 2022) training set and posed 103 basic understanding questions. We then evaluated various VLMs on these questions. The results, presented in the Appendix, show that performance dropped by over 40% compared to ChartQA scores, even for questions on training set charts. This highlights two points: first, current VLMs are capable of answering some chart-reasoning questions, but they lack a comprehensive understanding of charts. Second, the ChartQA dataset allows models to correctly answer questions without a comprehensive understanding of the charts, leading to an overestimation of the capabilities of current models (Wang et al. 2024).
@@
Firstly, the lack of high-quality chart training data is a major reason why current models lack robust chart understanding capabilities. Existing methods (Masry et al. 2022, 2023; Meng et al. 2024) for collecting chart training data fall into two categories: manual annotation and automatic synthesis. Manually annotated data have real-world chart appearances but suffer from coarse granularity and high human costs. Automatically synthesized data offer fine-grained annotations but lack real-world diversity, leading to poor model robustness. Thus, constructing chart datasets is a challenging balance between cost and quality, resulting in a scarcity of high-quality chart training data.
@@
  • We conducted extensive experiments on the EvoChart method and EvoChart-QA. Results demonstrate that the EvoChart method significantly outperforms other data synthesis methods, and we also deeply analyze the performance of various VLMs on EvoChart-QA.
@@ 2.1 Chart Question Answering Datasets
Since FigureQA (Kahou et al. 2018) pioneered the CQA task, numerous datasets for chart question answering have emerged. Synthetic datasets, such as DVQA (Kafle et al. 2018), PlotQA (Methani et al. 2020), RealCQA (Ahmed et al. 2023), ChartX (Xia et al. 2024) and UniChart (Masry et al. 2023). This datasets utilize synthetically generated charts or templat-base questions. Generate datasets such as ChartSFT (Meng et al. 2024), utilize a mixture of GPT-4 (OpenAI et al. 2024)-generated charts and questions. Mixed datasets as ChartQA (Masry et al. 2022) and Charxiv (Wang et al. 2024), the former is a dataset compiled semi-manually with the assistance of templates, while the latter is template-based and requires evaluation by GPT-4o. In contrast, EvoChart-QA focus on real-world scenarios and employ an automated evaluation method that does not necessitate the utilization of GPT-4.
@@ 2.2 Visual Language Models on CQA
VLMs are language models with visual understanding capabilities, and they have numerous applications in CQA tasks. Small VLMs like ChartReader (Cheng, Dai, and Hauptmann 2023), MatCha (Liu et al. 2023a), ScreenAI (Baechler et al. 2024) and UniChart (Masry et al. 2023) have shown superior performance on tasks like PlotQA and DVQA, highlighting the potential of VLMs in CQA. ChartLlama (Han et al. 2023) was a milestone, being the first to apply LLaVa1.5 (Liu et al. 2024) to CQA tasks and achieving impressive performance. Subsequently, works such as ChartPaLI (Carbune et al. 2024), ChartInstruct (Masry et al. 2024a), ChartAst-D (Meng et al. 2024), and TinyChart (Zhang et al. 2024) delved into the multimodal alignment and CQA reasoning aspects of VLMs in CQA, achieving remarkable performance. Recently, open-source general VLMs such as Phi3-Vision (Abdin et al. 2024) and Intern-VL2.0 (Chen et al. 2023b), through large-scale training, have achieved state-of-the-art performance on the ChartQA dataset.
@@ 2.3 Self-Training Approach
With the increasing capabilities of language models, numerous researchers have begun to explore the potential of leveraging language models for self-training. GPT3Mix (Yoo et al. 2021) proved that large language model augmentation of textual corpora is very effective. Further, ReST (Gülçehre et al. 2023) achieves cost-effective and efficient human preference alignment through a dual-loop self-training approach. Dennis et al. (Ulmer et al. 2024) obtained new data from the self-talk of multi-role-playing LLM Agents by adding a filtering check mechanism, realizing efficient self-training. Recently, Xu et al. (Xu et al. 2024) proposed ENVISIONS, which uses a neural-symbolic self-training approach to significantly improve mathematical and logical reasoning abilities without relying on external stronger models or evaluation tools. Inspired by their work, our proposed EvoChart focuses on a scalable self-training process.
@@ 3.1 Compositional Chart Generation
Compositional chart generation aims to produce high-quality and diverse charts at minimal cost, serving as the core component of EvoChart’s construction. Previous approaches (Han et al. 2023; Meng et al. 2024; Zhang et al. 2024) either rely solely on GPT-4 for chart generation result in limited diversity and high costs, or using plotting libraries for random generation often leads to unrealistic themes and styles. To achieve continuous, low-cost, and diverse chart creation, we propose a two-step generation strategy within the compositional chart generation process:
1) Chart Code Seed Generation: This step serves as the initial phase of chart generation, aiming to produce fundamental chart code containing elements that are difficult to achieve via random processes. These elements include chart themes and appropriate units for the x- and y-axes. Notably, this step is executed only once during the entire chart dataset construction process. As (Xu et al. 2023) demonstrated the feasibility of large language models for code generation, we employ sophisticated prompt engineering techniques to guide GPT-4 in generating over 25k real-world chart code seeds. Our GPT-4 prompts are based on the following key aspects:
@@
Other Configurations: A multitude of detailed configurations influence chart appearance, including line width, numeric label (position), line style (solid or dashed), bar stacking, axis visibility, font size, font type, and more. We introduce randomness into these configurations through a range of selectable options, and we employ ECharts (Li et al. 2018) for rendering charts.
@@ 3.2 Chart Evaluation and Refinement
3.3 QA-pairs Generation and Training
QA-pairs Generation and Training aims to generate chart-based question-answer pairs, incorporating these data into the EvoChart corpus and training the EvoChart model for the next stage. We generate question-answer pairs using various question templates. Notably, we focus on basic chart understanding, the templates specifically focus on the alignment of visual-text-value information in charts (e.g., extracting values through visual information, extracting visual information through text). Additionally, we generate rich question-CoT pairs using composable vCoTs (Rose et al. 2024) and distinguish Direct from vCoT using Instruct. Since vCoTs solely serve to enhance model comprehension, we only mix in 20% of vCoT data in the training data. We have established 198 question templates with corresponding answers including Direct and over 500 vCoT templates. We generated 1.6M QA-pairs during the training in 3 stages. A detailed information can be found in the Appendix.
@@ 4 EvoChart-QA Benchmark
4.1 Chart Selection
To enable EvoChart-QA to emulate real-world chart understanding scenarios, all charts in our dataset are carefully selected by human experts. Specifically, we crawled 1,000 charts from 140 different websites. Human experts then filtered out images with ambiguous meanings or damages, resulting in a final dataset of 625 valid images. These images include line charts, bar charts, pie charts, and scatter plots. Examples and sources of all images are detailed in the Appendix.
4.2 Question Construction
We focus on chart basic understanding questions. Following the definitions of prior researchers (Kafle et al. 2020; Methani et al. 2020), we concentrate on data and structural retrieval questions. Specifically, we categorize the problems into two types during manual construction: Direct Retrieval and Complex Retrieval. Direct Retrieval questions focus on understanding the image and directly extracting its content, while Complex Retrieval questions emphasize performing multiple visual reasoning steps on the chart.
1) Direct Retrieval. Direct Retrieval aims to directly extract elements from a chart based on the question. To comprehensively assess the model’s ability to extract various elements from charts, we categorize chart elements into three types: label, value, and visual. Label elements refer to textual content in the chart, such as chart title, axis labels, etc. Value elements refer to data values conveyed by the chart, which are displayed or implicitly provided based on the chart author’s intention. Visual elements refer to all visual descriptions in the chart, such as line color, largest segment, etc. For example: “What is the value of the green dashed line in 2015?” Although Direct Retrieval involves only extracting elements from charts, it remains a challenging task. On the one hand, real-world charts often exhibit non-standard variations. For example, they may include rich text such as numerous logos inserted in the image, or they may combine multiple chart forms within a single chart to facilitate expressions. On the other hand, questions posed by real-world users may contain ambiguous expressions. For example, “Which country’s total GDP is represented by the bar in the lightest shade of blue?” This visual description is ambiguous, but it is a clearly question for human observers.
@@ 5.1 Setup
Datasets. To comprehensively evaluate the effectiveness of EvoChart, we chose to test it on both ChartQA and EvoChart-QA. ChartQA (Masry et al. 2022) is a dataset with two subsets: “Augment”, which is machine-generated, and “Human”, which is manually curated. “Augment” focuses on element extraction tasks within machine-synthesized images, while “Human” emphasizes complex numerical and logical reasoning tasks in real-world charts. EvoChart-QA is a novel real-world benchmark that we proposed.
Models. We conducted extensive evaluations on both open-source and proprietary models. For open-source models, we tested Phi3-Vision-4B (Abdin et al. 2024), QwenVL-Chat-7B (Bai et al. 2023), LlaVa1.6-Vicuna-7B (Liu et al. 2024), Intern-VL-2.0-8B (Chen et al. 2023b), Llama3-Llava-Next-8B (Li et al. 2024), CogVLM2-19B (Wang et al. 2023), LlaVa1.6-YI-34B, Intern-VL-2.0-40B, ChartLlama-13B (Han et al. 2023), ChartAst-S-13B (Meng et al. 2024), ChartIns-Llama2-7B (Masry et al. 2024a), ChartIns-FlanT5-3B, ChartGemma-2B (Masry et al. 2024b), and TinyChart-3B (Zhang et al. 2024). For proprietary models, we tested Gemini-1.5-Flash (Team et al. 2024), Gemini-1.5-Pro, Qwen-VL-Plus, Qwen-VL-Max, GPT-4-turbo (OpenAI et al. 2024), and GPT-4o. For all models, we employed a zero-shot approach. The specific configurations of all models are provided in the Appendix.
Settings. In EvoChart method, we utilize Phi3-Vision (Abdin et al. 2024) as the initialization model. We conducted a 3-Stage data synthesis and training process, with each Stage undergoing 1 Epoch of SFT with a learning rate of 2e-5 and using cosine learning rate scheduler. All experiments were completed on 4 NVIDIA A800 80G GPUs.
@@ 5.2 Experimental Results
Tables 2 and 3 present the performance of EvoChart and other open-source or proprietary VLMs on EvoChart-QA and ChartQA. We elaborate on the experimental results in two aspects: EvoChart-QA and EvoChart.
@@
2) All models demonstrate significantly weaker performance on Complex Retrieval compared to Direct Retrieval, indicating that reasoning over visual information poses a substantially greater challenge than direct extraction of information from charts. Furthermore, nearly all models exhibit lower accuracy on Pie and Scatter chart types compared to their average performance. This suggests that Pie and Scatter charts pose a greater challenge.
3) Open-source general-purpose models exhibit comparable chart comprehension abilities to proprietary models. This suggests that models pretrained on large-scale chart data possess strong generalization capabilities. However, while chart expert models fine-tuned on specific domains achieve impressive scores on the ChartQA dataset, their performance degrades significantly to below 30% when confronted with the entirely OOD EvoChart-QA dataset.
@@
1) Although the EvoChart model is trained on synthetic data, it achieves SoTA performance on the entirely real-world benchmark EvoChart-QA and exhibits competitive performance on the chart reasoning task ChartQA. This validates the strong generalization ability of the EvoChart.
@@ 5.3 Analysis
EvoChart Ablation Study. We conducted comprehensive ablation studies on EvoChart, and the results are summarized in Table 4. We trained and generated EvoChart for 1 to 3 stages. Meanwhile, to verify the effectiveness of Chart Evaluation and Refinement, we set up EvoChart without refinement and trained it for 3 stages, denoted as “w/o refine stage-3.” We observed the following:
1) As the number of EvoChart method Stages increases, the scale of the EvoChart-Dataset expands, and the performance of the EvoChart steadily improves. This highlights EvoChart’s effectiveness as a self-training approach.
@@
6 Conclusion
In this paper, we introduce EvoChart and EvoChart-QA: a novel approach for enhancing chart comprehension capabilities through self-training and iterative synthetic data generation, and a meticulously crafted real-world chart comprehension benchmark. We aim to provide a new avenue for real-world chart understanding through EvoChart and EvoChart-QA. Through extensive experimentation, we expose the limitations of existing VLMs in chart comprehension and validate the effectiveness of our EvoChart method across multiple datasets. In the future, we will further explore human-free methods in chart comprehension.